# Shape-specific fluctuations of an active colloidal interface

### v1

@@ Abstract
Motivated by a recently synthesized class of active interfaces formed by linked self-propelled colloids, we investigate the dynamics and fluctuations of a phoretically (chemically) interacting active interface with roto-translational coupling. We enumerate all steady-state shapes of the interface across parameter space and identify a regime where the interface acquires a finite curvature, leading to a characteristic “C-shaped” topology, along with persistent self-propulsion. In this phase, the interface height fluctuations obey Family–Vicsek scaling but with novel exponents: a dynamic exponent $z_{h}\approx 0.6$, a roughness exponent $\alpha_{h}\approx 0.9$ and a super-ballistic growth exponent $\beta_{h}\approx 1.5$. In contrast, the orientational fluctuations of the colloidal monomers exhibit a negative roughness exponent, reflecting a surprising smoothness law, where steady-state fluctuations diminish with increasing system size. Together, these findings reveal a unique non-equilibrium universality class associated with self-propelled interfaces of non-standard shape.
@@ I Introduction
Out-of-equilibrium agents with suitably chosen interactions are known to exhibit emergent collective order, most prominently in the form of global polar order  [1, 2]. Such order can arise from alignment rules [2, 3], topological interactions [4], long-ranged attraction or repulsion [5, 6], and/or other generic behavioral couplings [7]. In parallel, a distinct body of work in statistical physics has established a deep understanding of fluctuating interfaces, where the statistical properties of the interface height obey universal scaling laws [8, 9, 10, 11]. A cornerstone result in this field is the Family–Vicsek (FV) scaling law, which relates the roughness of the interface to the system size and time through universal exponents [12, 13]. These exponents are known for certain solvable models, such as the Edwards–Wilkinson (EW) and Kardar–Parisi–Zhang (KPZ) equations [8, 9, 14], while numerical and experimental studies have revealed a wealth of deviations, suggesting novel universality classes [15, 16, 17, 18]. Bridging these two domains, the study of active interfaces, i.e., interfaces driven out of equilibrium by internal activity, is a subject of recent interest  [19, 20, 21, 22, 23], raising new questions about how activity reshapes interfacial fluctuations and scaling behavior.
@@
In this article, we study a polar colloidal chain as a non-equilibrium interface in $1+1$ dimensions. This is inspired by recent experiments that have been able to synthesize an autonomous polar chain via chemical self-interactions [24], as opposed to other propulsion mechanisms via external actuation [25, 26, 27]. We study the dynamical steady-states of this system and show that within a specific parameter regime the chain ballistically propels with a deterministic “C-shape” (reported previously in [24, 28]). As opposed to conventional growing (circular) interfaces, the non-equilibrium interface we study here displays the additional phenomena (in addition to being propelled) of curvature acquisition during its dynamics, attained via the inter-monomeric phoretic interactions, thus adding additional time scales to the conventional early time deterministic plus late time diffusive behaviour. The polar nature of the interface (and hence the reason it breaks translational symmetry and propels in a given direction) enables calculations of height fluctuations of the monomers about a mean height [10] defined across the interface.
@@
We report that the interface height fluctuations of this chain obey FV scaling, characterized by a dynamic ($z_{h}\approx 0.6$) and a roughness ($\alpha_{h}\approx 0.9$) exponent. Correspondingly, the fluctuations preceding the steady state are captured by a super-ballistic growth exponent $\beta_{h}\approx 1.5$. The steady-state orientational fluctuations of the colloidal monomers, in contrast, decrease with the chain length, exhibiting a negative roughness (i.e. “smoothness”) exponent ($\alpha_{\theta}\approx-0.5$). Moreover, we report a distinct set of scaling exponents characterizing the locally flat regime, defined by monomer-level averages restricted to flat portions of the chain and, independently, by taking the infinite-chain limit. Given that the shape of the propelling interface does not fall under conventionally studied circular or flat interfaces [29], these exponents thus constitute a unique non-equilibrium signature of an interface with a stereotypic “C-shape” topology.
@@
The paper is structured as follows. In Sect. II, we introduce the model and enumerate the relevant length and time scales of our model. In Sec. III , we study the dynamical regimes of the model as a function of selected dimensionless numbers and obtain a phase diagram [Fig. 1]. We report the results for the scalings of height and orientation fluctuations in Sec. IV. We finally discuss the significance of our findings and future directions in Sec. V.
Figure 1: Phase diagram in $\Lambda$-$\tilde{\Lambda}$ plane. The log curvature has been used to delineate the phases. The phase diagram has been drawn with for a chain with $N=256$ number of monomers. Representative images for each phase is shown along with the marker key for a smaller chain for clarity. See text for details.
@@ II Model
We model the $i$th active droplet as a colloid particle centered at $\mathbf{r}_{i}=(x_{i},y_{i})$. The particle is confined to move in two-dimensions. It self-propels with a speed $v_{s}$, along the directions $\mathbf{e}_{i}=(\cos\theta_{i},\,\sin\theta_{i})$. The direction of the particle, given by the angle $\theta$, changes due to coupling to a phoretic field $c({\bf r},t)$. The position and orientation of the $i$th particle are determined by the following evolution equations:
@@
In the above equations, $\mu$ is mobility, $D_{t}$ and $D_{r}$, respectively, are translational and rotational diffusion constants of the particle, while $\bm{\xi}^{t}$ and $\bm{\xi}^{r}$ are white noises with zero mean and unit variances. The constant $\chi_{r}$ is taken to be positive so that the particles rotate away from each other (chemo-repulsive). The force on the $i$th particle is given as:
$\mathbf{F}_{i}^{b}=-\frac{\partial U}{\partial\mathbf{r}_{i}},\qquad U=\sum^{N-1}_{i=1}\mathcal{U}^{b}(\mathbf{r}_{i},\mathbf{r}_{j}).$
Here, $\mathcal{U}^{b}(\mathbf{r}_{i},\mathbf{r}_{j})=k\left(r_{ij}-r_{0}\right)^{2}$ is a harmonic potential of stiffness $k$ and natural length $r_{0}$ which holds the chain together, while $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$. The chemical interactions between the monomers of the chain are contained in $\mathbf{J}_{i}=-\partial c/\partial\mathbf{r}_{i}$, where $c$ is the concentration of the phoretic field. It is worthwhile to note that the positional and orientational dynamics are coupled in our model. This roto-translational coupling leads to rich phenomenology of our model, as we describe below.
In the steady-state, the solution of the concentration profile $c(\bm{r})$ follows from the equation: $D_{c}\nabla^{2}c(\boldsymbol{r},t)+\sum_{i=1}^{N}c_{0}\,\delta(\boldsymbol{r}-\boldsymbol{r}_{i})=0$, where $D_{c}$ is the diffusion coefficient of the filled micelles and $c_{0}$ is emission constant of the micelles. This gives the expression for the current:
@@
This current is thus interpreted as that which is instantaneously deposited on the centers of the colloids. For the entire paper (unless specified otherwise) we use the following initial conditions:
@@
where $b$ is the radius of each colloid. Note that $r_{0}=2b$. The above model has been studied recently [24, 28]; a detailed phase diagram of the model was obtained in [28] in terms of the dimensionless group of the system. Here, we focus on the fluctuations in the so-called “C-shape” phase, where the chain spontaneously acquires a stereotypic shape resembling the alphabet C, and propels in a direction normal to its tangent [24, 28]. We note that the “C-shape” topology we refer to is specific to systems in which roto-translational coupling in the dynamics induces the steady-state shape. The system we study being “dry” with such a coupling, is distinct from other similar topologies seen in (for instance) the physics of sedimenting filaments [30]. In the limit of $N\rightarrow\infty$ it corresponds to an interface which is moving. A schematic of this is shown in Fig. 2. We focus on the fluctuations of that moving interface in this paper.
All numerical results in this paper are generated with an explicit Euler-Maruyama integrator with time step $dt=0.01$ and simulation time $T=7.2\times 10^{5}$. For all simulations here (unless stated otherwise) $b=1$, $c_{0}=1$, $D_{c}=1$, $\hat{\mu}=\mu k_{b}=91.6s^{-1}$, and $v_{s}=1$. Wherever realization average is performed, we use $50$ realizations, with the exception of $10$ used in Appendix .7.
@@ III Dynamical regimes
We first define some typical time scales, ratios of which determine the relevant dimensionless numbers needed to understand the phase diagram (Fig. 1):
@@
We now present the dynamics regimes of the model described above in terms of the two dimensionless numbers. In Fig. 1, we present a phase diagram in the plane of $\Lambda$ and $\tilde{\Lambda}$. We find that the C-shape is sustained for a selected range of dimensionless numbers $\Lambda$ and $\tilde{\Lambda}$; the details of this phase have already been discussed in [28]. For regions $\tilde{\Lambda}<10^{0}$, we find a disordered phase with effectively zero positional order in the chain. Using the exact simulation values, we may define a length scale $\frac{D_{t}}{v_{s}}$ which would correspond to $\frac{D_{t}}{v_{s}}>b$, thus we conclude that long-wavelength fluctuations (that is, greater than the monomer size) are not supported by our colloidal chain. Focusing in the regime $\tilde{\Lambda}>10^{1}$ and $\Lambda<10^{-3}$ corresponds to the case where deterministic rotations are instantaneous across the system ($\tau_{f}<<\tau$), such that the chemical interactions are instantaneously correlated. Again, by constructing a typical length scale one finds in this regime that the deterministic interactions $\gtrsim 10^{3}b$, thus correlations much larger than the system size. The dynamics of the chain displays “frustrated” behavior, not attaining any non-trivial spatial structure as there is not enough time to respond to instantaneous chemical gradients. Such a phase was not reported in [28] due to the existence of chemical trails that break the forward-backward symmetry; this is also discussed in Appendix .7. The limit of $\Lambda>10^{3},\tilde{\Lambda}>10^{0}$ corresponds to the situation when there are insufficient rotations and the chain remains rigid within the simulation time scale [28]. This instead corresponds to the region of extremely short-wavelength fluctuations and small deterministic correlation lengths. We thus conclude that the C-shape thus corresponds to regimes where the correlation lengths $\frac{\chi_{r}}{b^{3}}$ are of the order of the chain length, which is in addition stable to short-wavelength fluctuations (of typical distance less that one monomer radius). It is this dynamical steady-state (DSS) phase that we study in the rest of this paper.
@@
We also note that, within the C-shape phase, its existence as a DSS depends on the choice of initial conditions of Eq. (4). Deviating from these does not render a universality of the DSS, unless one includes the existence of chemical trails in the model, as discussed in [24, 28]. This scenario is further discussed in Section .7. If one deviates from the purely chemo-repulsive scenario considered here, the presence of trails are also necessary for other non-trivial DSS, such as the swimming/undulating state. However, various structural SS (e.g. designer crystallites that break handedness symmetry) is possible via the instantaneous chemical deposition model considered here [28].

### v2

@@ Abstract
Motivated by a recently synthesizable class of active interfaces formed by linked self-propelled colloids, we investigate the dynamics and fluctuations of a phoretically (chemically) interacting active interface with roto-translational coupling. We enumerate all steady-state shapes of the interface across parameter space and identify a regime where the interface acquires a finite curvature, leading to a characteristic “C-shaped” topology, along with persistent self-propulsion. In this phase, the interface height fluctuations obey Family–Vicsek scaling but with novel exponents: a dynamic exponent $z_{h}\approx 0.5$, a roughness exponent $\alpha_{h}\approx 0.9$ and a super-ballistic growth exponent $\beta_{h}\approx 1.7$. In contrast, the orientational fluctuations of the colloidal monomers exhibit a negative roughness exponent, reflecting a surprising smoothness law, where steady-state fluctuations diminish with increasing system size. Together, these findings point towards a unique non-equilibrium universality class associated with self-propelled interfaces of non-standard shape.
@@ I Introduction
Out-of-equilibrium agents with suitably chosen interactions are known to exhibit emergent collective order, most prominently in the form of a global polar order  [1, 2]. Such order can arise from alignment rules [2, 3], topological interactions [4], long-ranged attraction or repulsion [5, 6], and/or other generic behavioral couplings [7]. In parallel, a distinct body of work in statistical physics has established a deep understanding of fluctuating interfaces, where the statistical properties of the interface height obey universal scaling laws [8, 9, 10, 11]. A cornerstone result in this field is the Family–Vicsek (FV) scaling law, which relates the roughness of the interface to the system size and time through universal exponents [12, 13]. These exponents are known for certain solvable models, such as the Edwards–Wilkinson (EW) and Kardar–Parisi–Zhang (KPZ) equations [8, 9, 14], while numerical and experimental studies have revealed a wealth of deviations, suggesting novel universality classes [15, 16, 17, 18]. Bridging these two domains, the study of active interfaces, i.e., interfaces driven out of equilibrium by internal activity, is a subject of recent interest  [19, 20, 21, 22, 23, 24], raising new questions about how activity reshapes interfacial fluctuations and scaling behavior.
@@
In this article, we study a polar colloidal chain as a non-equilibrium interface in $1+1$ dimensions. This is inspired by recent experiments that have been able to synthesize an autophoretic polar chain via chemical self-interactions [25], as opposed to other propulsion mechanisms via external actuation [26, 27, 28]. We study the dynamical steady-states of this system and show that within a specific parameter regime the chain ballistically propels with a deterministic “C-shape”[25, 29]. As opposed to conventional growing (circular) interfaces, the non-equilibrium interface we study here displays the additional phenomena (in addition to being propelled) of curvature acquisition during its dynamics, attained via the inter-monomeric phoretic interactions, thus adding additional time scales to the conventional early time deterministic plus late time diffusive behaviour [30]. The polar nature of the interface of this topology (and hence the reason it breaks translational symmetry and propels in a given direction) enables calculations of height fluctuations of the monomers about a mean height defined across the interface.
@@
We report that the interface-height fluctuations of this chain obey FV scaling, characterized by a set of dynamic ($z_{h}\approx 0.5$) and roughness ($\alpha_{h}\approx 0.9$) exponents. Correspondingly, the fluctuations preceding the steady state are captured by a super-ballistic growth exponent $\beta_{h}\approx 1.7$. The steady-state orientational fluctuations of the colloidal monomers in the C-shape phase, in contrast, decrease with the chain length, exhibiting a negative roughness (i.e. “smoothness”) exponent ($\alpha_{\theta}\approx-0.5$). Moreover, we report a distinct set of scaling exponents characterizing the locally flat regime, defined by monomer-level averages restricted to flat portions of the chain and, independently, by taking the infinite-chain limit. Given that the shape of the propelling interface does not fall under conventionally studied circular or flat interfaces [31], these exponents thus constitute a unique non-equilibrium signature of an interface with a stereotypic “C-shape” topology.
@@
The paper is structured as follows. In Sect. II, we introduce the model and enumerate the relevant length and time scales of our model. In Sec. III , we study the dynamical regimes of the model as functions of selected dimensionless numbers and obtain a phase diagram [Fig. 1]. The main results of this paper are presented in Sec. IV, where we discuss in detail the height and orientational fluctuations of this system, in addition to the effect of finite rotational noise. We finally discuss the significance of our findings and future directions in Sec. VI.
[…]
@@ II Model
We consider an active polymer consisting of chemically interacting active particles as monomers. The $i$th active particle - centered at $\mathbf{r}_{i}=(x_{i},y_{i})$ - is confined to move in two-dimensions. It self-propels with a speed $v_{s}$, along the directions $\mathbf{e}_{i}=(\cos\theta_{i},\,\sin\theta_{i})$. The position and orientation of the $i$th particle are determined by the following evolution equations:
@@
In the above equations, $\mu$ is mobility, $D_{t}$ and $D_{r}$, respectively, are translational and rotational diffusion constants of the particle, while $\bm{\xi}^{t}$ and $\bm{\xi}^{r}$ are white noises with zero mean and unit variances. The constant $\chi_{r}$ is taken to be positive so that the particles rotate away from each other (chemo-repulsive). The force on the $i$th particle is given as: $\mathbf{F}_{i}^{b}=-{\partial U}/{\partial\mathbf{r}_{i}}$, while we have: $U=\sum^{N-1}_{i=1}\mathcal{U}^{b}(\mathbf{r}_{i},\mathbf{r}_{i+1})$. Here, $\mathcal{U}^{b}(\mathbf{r}_{i},\mathbf{r}_{j})=k\left(r_{ij}-r_{0}\right)^{2}$ is a harmonic potential of stiffness $k$ and natural length $r_{0}$ which holds the chain together, while $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$.
The orientation of the particles, given by the angle $\theta_{i}$, changes due to coupling to a chemical (phoretic) field $c({\bf r},t)$. The chemical interactions between the monomers of the chain are contained in $\mathbf{J}_{i}=-\partial c/\partial\mathbf{r}_{i}$, where $c$ is the concentration of the phoretic field. It is worthwhile to note that the positional and orientational dynamics are coupled in our model. This roto-translational coupling leads to rich phenomenology of our model, as we describe below. In the steady-state, the solution of the concentration profile $c(\bm{r})$ follows from the equation: $D_{c}\nabla^{2}c(\boldsymbol{r},t)+\sum_{i=1}^{N}c_{0}\,\delta(\boldsymbol{r}-\boldsymbol{r}_{i})=0$, where $D_{c}$ is the diffusion coefficient of the filled micelles and $c_{0}$ is emission constant of the micelles. This gives the expression for the current:
@@
This current is thus interpreted as that which is instantaneously deposited on the centers of the colloids.
Variants of the above model have been studied recently [25, 29]; a detailed phase diagram of the model was obtained in [29] in terms of the dimensionless group of the system. Here, we focus on the fluctuations in the so-called “C-shape” phase, where the chain spontaneously acquires a stereotypic shape resembling the alphabet C, and propels in a direction normal to its tangent [25, 29]. We note that the “C-shape” topology we refer to is specific to systems in which roto-translational coupling in the dynamics induces the steady-state shape. The system we study being “dry” with such a coupling, is distinct from other similar topologies seen in (for instance) the physics of sedimenting filaments [32]. In the limit of $N\rightarrow\infty$ it corresponds to an interface which is moving. A schematic of this is shown in Fig. 1D. We focus on the fluctuations of that moving interface in this paper.
For the entire paper (unless specified otherwise), we use the following initial conditions:
@@
where $b$ is the radius of each colloid. Note that $r_{0}=2b$. All numerical results in this paper are generated with an explicit Euler-Maruyama integrator with time step $dt=0.01$ and simulation time $T=7.2\times 10^{5}$. For all simulations here (unless stated otherwise) $b=1$, $c_{0}=1$, $D_{c}=1$, $\hat{\mu}=\mu k_{b}=91.6s^{-1}$, and $v_{s}=1$. $D_{r}=0$ for all sections, with the exception of Sec. V.3 where values are explicitly stated. Wherever realization average is performed, we use $50$ realizations, with the exception of $10$ used in Appendix .1.
@@ III Dynamical regimes
We first define some typical time scales, ratios of which determine the relevant dimensionless numbers needed to understand the non-equilibrium phase diagram (Fig. 1):
@@
It is informative to study the model described above in terms of the two dimensionless numbers. In Fig. 1, we present a phase diagram in the plane of $\Lambda$ and $\tilde{\Lambda}$. We find that the C-shape is sustained for a selected range of dimensionless numbers $\Lambda$ and $\tilde{\Lambda}$ [25, 29]. Let us construct in addition a fluctuating length scale $l_{f}=\frac{D_{t}}{v_{s}}$ and a correlation length scale $l_{C}=\frac{\chi_{r}}{bv_{s}}$ to aid with the analysis. First, in the limit of $\Lambda\gtrsim 10^{2},\tilde{\Lambda}\gtrsim 10^{1}$, there exists insufficient rotations and the chain remains flat within the simulation time scale [29]. This instead corresponds to the region of extremely short-wavelength fluctuations ($l_{f}<b$) and small deterministic correlation lengths ($l_{C}<10^{-3}b$). This straight-chain phase is labelled as Phase I (see Fig. 1(b) top left). Next, for regions $\tilde{\Lambda}\lesssim 10^{0}$, we find a disordered phase with effectively zero positional order in the chain. Here we find that $l_{f}>b$, thus we conclude that long-wavelength fluctuations of greater than the monomer size are not supported by our colloidal chain. This disordered phase is labelled as Phase II (Fig. 1(b) top right). Further, focusing in the regime $\tilde{\Lambda}>10^{1}$ and $\Lambda<10^{-3}$ corresponds to the case where deterministic rotations are instantaneous across the system ($\tau_{f}<<\tau$), such that the chemical interactions are instantaneously correlated. Here one has that $l_{C}\gtrsim 10^{3}b$, thus correlations much larger than the system size. The dynamics of the chain displays “frustrated” behavior, not attaining any non-trivial spatial structure as there is not enough time to respond to instantaneous chemical gradients. This frustrated phase is labelled as Phase III (Fig. 1(c) bottom left). The remaining region corresponds to that of the C-shape (Phase IV). We thus conclude that the C-shape thus corresponds to regimes where the correlation lengths $l_{C}$ are of the order of the chain length, which is in addition stable to short-wavelength fluctuations (of typical distance less that one monomer radius). It is this dynamical steady-state (DSS) phase that we study in the rest of this paper.
@@
We also note that, within the C-shape phase, its existence as a DSS depends on the choice of initial conditions of Eq. (4). Deviating from these does not render a universality of the DSS, unless one includes the existence of chemical trails in the model [25, 29]. This scenario is further discussed in Appendix .1. If one deviates from the purely chemo-repulsive scenario considered here, the presence of trails are also necessary for other non-trivial DSS, such as the swimming/undulating state. However, various structural SS (e.g. designer crystallites that break handedness symmetry) is possible via the instantaneous chemical deposition model considered here, though one would have to significantly alter from the fully chemo-repulsive scenario, and incorporate mixed (anti-symmetric and/or non-reciprocal) interactions [29].
III.1 Characteristics of the C-shape
[…]
  1. 1.
    Continuously varying orientations along the chain
  2. 2.
    Finite curvature along chain (or the averaged local curvature)
  3. 3.
    Ballistically propelling in one direction
  4. 4.
    Positional and orientational symmetry along the body(y) axis
As explained in Appendix .7, there are various other DSS that meet some of the above requirements but not all. For instance the aforementioned Phase I DSS satisfies only (i), (iii) and (iv). The C-shape is unique in that it is the only DSS that meets requirements (i)-(iv), with its propulsion perpendicular to the direction of the body axis. It is in this specific phase in which the fluctuations are computed.
Section Quantity $\beta^{(1)}$ $\beta^{(2)}$ $\alpha$ $z=\alpha/\beta$ Figure Remarks
V.1 $W_{h}$   $0.25\pm 0.01\,\,$   $1.73\pm 0.01$   $0.90\pm 0.03\,\,$   $0.52\pm 0.02\,\,$ 2 Phase IV - positional
V.2 $W_{\theta}$   $1.00\pm 0.01\,\,$ NA $0.48\pm 0.02\,\,$ $0.48\pm 0.02\,\,$ 3 Phase IV - orientational
V.3 $W_{h}$   $0.25\pm 0.02\,\,$   $1.73\pm 0.03\,\,$   $0.81\pm 0.04\,\,$   $0.47\pm 0.03\,\,$ 4 Finite $\tau_{r}$
$W_{\theta}$   $1.00\pm 0.02\,\,$ NA $0.48\pm 0.02\,\,$ $0.48\pm 0.02\,\,$
V.4 $W_{h}$   $0.25\pm 0.02\,\,$ NA   $1.00\pm 0.01\,\,$   $4.00\pm 0.30\,\,$ 5 Locally flat - positional
Appendix .6 $W_{\theta}$   $1.00\pm 0.01\,\,$ NA   $1.02\pm 0.02\,\,$   $1.02\pm 0.02\,\,$ 8 Locally flat - orientational
Table 1: Table of exponents defined in equations (8-14) of section IV. Note that $z$ is defined such that $\beta^{(2)}$ is in the denominator for the first and third rows, whereas $\beta^{(1)}$ is in the denominator for the others (orientational and locally flat exponents). For the last two columns, $z$ is taken to be $z^{\prime}$ implicitly. For comparison, we note that for a one-dimensional interface: $\alpha^{\mathrm{EW}}=1/2$, $\beta^{\mathrm{EW}}=1/4$; $\alpha^{\mathrm{KPZ}}=1/2$, $\beta^{\mathrm{KPZ}}=1/3$ [34, 35].
III.2 C-shape selection
In Phase IV, as mentioned previously, the C-shape is always attained under the initial conditions of 4. We note that in such an autophoretic chain with chemical trails this DSS is the universal solution regardless of initial conditions [25, 29] (see Appendix .1)). However, it is instructive to ask to what extent the C-shape is attained under different initial conditions.
To this end, we apply a hysteretic protocol by varying $\Lambda$ and carrying on the final state of a given simulation to the initial conditions of the simulation at the next time step. The order parameter used to identify this DSS is the absolute value of the curvature, $|\kappa|$. The results of these are presented in Fig. 1(c). Two separate cases are presented, corresponding to decrease of $\Lambda$ starting from Phase I (the stiff chain; Fig. 1(c), top), and increase of $\Lambda$ starting from Phase IV (the C-shape; Fig. 1(c), bottom). We observe from the former that the stiff-to-C transition is irreversible. Thus, a C-shape DSS cannot lose its spatial curvature via a decrease in self-phoretic torques. From the latter we see that the C-to-stiff transition is hence not permitted; both forward and backward sweeps of the protocol do not change the overall curvature. We also note that we have separately varied $\tilde{\Lambda}$ to and from Phase IV (not shown here); the results of these indicate that selected random configurations of Phase II may indeed transition to IV, but again there is no clear reliable reversibility of transition to define a hysteresis loop. This irreversibility of the first transition implies that such a hysteresis curve and hence mapping to equilibrium classes of phase transitions is not well defined (indeed one may have to devise an alternate order parameter); we leave the exact characterization of the nature of the transition to and from the C-shape for future work.
We merely conclude this section by emphasizing that the C-shape is the attractor (global minimum) of our EOM (2) with initial conditions of (4). The subsequent results of this work are thus based on the assumption of instantaneous chemical depositions and the aforementioned initial conditions; we note that these may change upon incorporation of chemical trails. We display in Appendix .1 that identical growth exponents are obtained when chemical trails are present; where the C-shape is instead a universal attractor (independent of initial conditions) [29].